OpenStudy (broskishelleh):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)). -40x + 24 -40x + 31 -40x + 64 -40x + 70

OpenStudy (broskishelleh):

@nikato

OpenStudy (anonymous):

Hey, so for these kinds of question you're basically substituting g(x) into f(x) where ever the x is. So, \[f(g(x)) \implies -4(10x-6)+7\] now evaluate

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ f(x) = -4x + 7\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ g(x) = 10x - 6\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ f(g(x)) = -4(10x - 6) + 7\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ f(g(x)) = -40x +24 + 7\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ f(g(x)) = -40x +31\LARGE\color{white}{ \rm │ }}\)

OpenStudy (broskishelleh):

Thanks @SolomonZelman Can I have One more question

OpenStudy (broskishelleh):

I have one more I mean

OpenStudy (solomonzelman):

sure

OpenStudy (solomonzelman):

shore

OpenStudy (broskishelleh):

Let f(x) = 2x - 6. Solve f^-1(x) when x = 2. -5 4 -2 7

OpenStudy (anonymous):

Let me let me!

OpenStudy (solomonzelman):

go ahead

OpenStudy (anonymous):

Ok when the notation f^-1(x) implies you're looking for the inverse, so first lets find the inverse... Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x) Once you have the inverse, sub 2 for x and evaluate :)!

OpenStudy (broskishelleh):

I don't get it

OpenStudy (anonymous):

\[f(x) = 2x-6\] we have to find the inverse, so following the steps I gave you, \[y=2x-6\] now switch the x's and y's around, \[x=2y-6\] now solve for y, \[y=\frac{ x+6 }{ 2 }\] \[\implies f ^{-1}(x) = \frac{ x+6 }{ 2 }\] now substitute 2 for x and solve.

OpenStudy (broskishelleh):

Okay so 8/2 =4

OpenStudy (broskishelleh):

So B

OpenStudy (broskishelleh):

One more please

OpenStudy (anonymous):

Yeah.

OpenStudy (broskishelleh):

Let f(x) = 24/3x -2 Find f(-2). -6 -3 3 6

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

If f(x) implies 24/3x-2 f(-2) implies ...?

OpenStudy (broskishelleh):

I believe it is 24/12

OpenStudy (anonymous):

Is it \[f(x) = \frac{ 24 }{ 3x-2 }~~~or~~~f(x) = \frac{ 24 }{ 3x }-2\]

OpenStudy (anonymous):

Either way just plug in -2 for wherever the x is.

OpenStudy (broskishelleh):

I am sorry I meant the 1st

OpenStudy (anonymous):

K well just plug in -2 for x

OpenStudy (broskishelleh):

3 x -2 = -6 - 2 = 8

OpenStudy (broskishelleh):

Ah so the answer is 3

OpenStudy (anonymous):

-3